OpenStudy (anonymous):

Write 738.01 as a mixed number.

OpenStudy (anonymous):

Well, the 738 is fine but the 0.1 can be changed into a fraction. That would create a mixed number.

OpenStudy (anonymous):

I dont know wat ur talking about

OpenStudy (snowcrystal):

738 and 1/100

OpenStudy (anonymous):

the answer is 738 1/100

OpenStudy (anonymous):

The answer should be explained and not only given, lol

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

so you can write it like? 7381/100

OpenStudy (anonymous):

\[738.01=\frac{ 73801 }{ 100 }=\frac{ 73800+1 }{ 100 }=\frac{ 73800 }{ 100 }+\frac{ 1 }{ 100 }\] \[=738+\frac{ 1 }{ 100 }=738\frac{ 1 }{ 100 }\]

OpenStudy (anonymous):

yo thanks

OpenStudy (anonymous):

yw